# Blockchain Economic Denial of Sustainability Attack: Exploiting Latency Optimization in Ethereum Transaction Forwarding

### v1

@@ Abstract
Strategies related to the blockchain concept of Extractable Value (MEV/BEV), such as arbitrage, front- or back-running create an economic incentive for network nodes to reduce latency, including minimizing transaction validation time—a core feature to secure blockchain networks. A modified node, that neglects to filter invalid transactions in the Ethereum P2P network, introduces novel attack vectors. In this work, we formalize and evaluate a Blockchain Economic Denial of Sustainability (EDoS) attack, which can cause financial losses in traffic costs for operators of modified nodes. We 1) mathematically define the attack model, 2) identify thousands of empirical instances of this similar attack in the wild, 3) empirically measure the model parameters from our two monitoring nodes, and 4) conduct attack simulations on the local network to compare its performance with existing Denial-of-Service attacks. We show that an attacker can amplify network traffic at modified nodes by a factor of 3,600, and cause economic damages 13,800 times greater than the amount needed to carry out the attack. Despite these risks, aggressive latency reduction may still be profitable enough to justify the existence of modified nodes. To assess this trade-off, we 1) simulate the transaction validation process in the local network and 2) empirically measure the latency reduction by deploying our modified node in the Ethereum testnet. We conclude with a cost-benefit analysis of skipping validation and provide mitigation strategies against this attack.
@@ 1 Introduction
The journey of transactions in blockchain P2P networks en route to their ultimate validation remains understudied. The specifics of handling pending transactions vary according to the client’s software version and its unique implementation. Under the concept of Extractable Value (MEV/BEV) [9, 44]—obtaining extra profits by reordering transactions, centralized business entities have the incentive to reduce latency to deliver pending transactions. Being the fastest in this space (as little as 1 millisecond) attracts users, bots, and validators who profit from arbitrage or front/back running [9, 45]. This scenario bears a resemblance to the traditional stock market where trading firms physically shortened the network paths to stock exchange matching engines and optimized networking configuration for quicker trade execution [33]. Recently, we notice certain modified nodes tailor their configurations, sidestep transaction validation processes to shorten latency, and introduce invalid transactions that are not supposed to persist in the network.
@@
Figure 1: Overview of Blockchain EDoS attack, illustrating how one invalid transaction amplifies in the network.
@@
We formalize the Blockchain Economic Denial of Sustainability (EDoS) attack, a superset of previously proposed DoS (Denial of Service) attacks [35, 54]. In Figure 1, the adversary sends an invalid transaction (red arrow) to the modified node(s) that collectively propagate an invalid transaction (blue arrows) to the rest of the network, causing increased outgoing network traffic, and as a result, financial losses for those modified nodes. This paper fully describes such potential attack opportunities, documents the prevalence of similar attacks in the wild, empirically/simulatively models the network damage, and quantifies the trade-offs associated with skipping validations.
@@
We first formally define the Blockchain EDoS attack and pinpoint the parameters that affect how an attacker’s invalid transactions can get amplified in the network. While considering the specifics of each software/version on transaction forwarding, our model establishes a framework to quantify the increase in both the amount and the economic cost of egress traffic at the modified nodes.
@@
We next find that some centralized services propagate a significantly higher number of invalid transactions by analyzing the set of pending transactions (called “txpool”) for each entity. We identify 2,591 similar attack instances across 345 Ethereum blockchain addresses that target these services; this substantiates the attack feasibility. We classify those instances based on the attacker’s observed strategies and estimate the attack cost, size, and intensity.
@@
To correctly estimate the overall impact of the attack and understand the topology of the network, we design two customized monitoring Geth nodes (the most popular Ethereum node client) to scan the activities on the P2P layer and record specific types of messages before they undergo processing, resulting 2.5 billion observations over 5 months. We formalize a new method for inferring the number of active peer connections, which serves as a cost-effective/ethical alternative to previously proposed methods [13, 10, 4, 34]. Furthermore, we show that 1.5% of nodes in our dataset exhibit lenient transaction validation. Many of those share an identical git commit, whose hash we do not find in any public software repository, implying those nodes come from the same entity.
Given the model we define and the parameters we obtain from our measurements, we show that the attacker can amplify outgoing network traffic at modified nodes by a factor of at least 3,600, and cause economic damages 13,800 times greater than the amount needed to carry out the attack.
@@
In addition, we conduct attack simulations on the local P2P network and confirm that our EDoS attack can evict as many honest transactions (from both the txpool/the block) as the existing DoS attacks but at significantly lower costs.
@@
Furthermore, we experiment to quantify the benefits of skipping transaction validation, specifically the amount of time saved for each validation process. We first fork our archive node (the node stores all the historical states of the blockchain) to simulate/dissect the process of Geth’s transaction validation in the locally controlled environment. We find that checking account status (nonce/balance) is the most time-consuming (86%). Moreover, we empirically measure the latency reduction by deploying the modified/regular nodes in the Ethereum testnet for 14 weeks.
@@
  • We formally define the Blockchain EDoS attack and show feasibility based on similar attacks in the wild.
  • By designing and deploying our custom monitoring nodes, we infer the Ethereum P2P network topology and calculate the attack amplification factors.
@@ 2.1 Blockchain and P2P network
Beyond protocol-level verifications, each software incorporates supplementary checks to mitigate DoS (Denial-of-Service) attacks. For example, Geth employs the “gas bump rule,” requiring the gas fee for a transaction with the same sender and nonce to be higher than the original transaction with a specified increase (e.g., 10% in Geth). The node should not accept replicated transactions with the same address, nonce, and gas price. Li et al. [34, §5.1] or Yaish et al.  [54, §4.1] provide detailed explanations.
@@
When users fetch the blockchain data or initiate a transaction, they typically interact with a P2P node through Remote Procedure Call (RPC) requests, and nodes execute these requests. Various publicly accessible RPC endpoints (e.g., Infura) allow users (rate limited) free access.
Txpool view providers (such as bloXroute, Eden, Chainbound) offer a “better” view of txpool in terms of latency. For example, bloXroute employs a Blockchain Distribution Network (BDN) infrastructure to improve blockchain scalability [31, 3]. The globally distributed nodes index transactions to reduce the size, internally propagate without validations (“cut-through routing”), and employ dynamic routing to overcome network overhead.
@@ 2.2 Security in P2P network
In this paper, we introduce an Economic Denial of Sustainability (EDoS) attack in blockchain P2P networks. Hoff [28] first presented the EDoS concept in a 2008 blog post. EDoS, akin to a traditional Denial of Service (DoS) attack, focuses on inflicting financial losses by targeting the victim’s traffic usage [52]. Attackers exploit the knowledge of a victim server hosted on a cloud service by manipulating traffic usage to inflate bills, making the victim’s operation economically unsustainable. The same issue applies to blockchain P2P networks where many nodes have public IP addresses, and use cloud platforms such as AWS.
@@ 3.2 Private monitoring node data
To effectively profile each node within a network, we gather peer information from our local Ethereum nodes. This is done by subscribing to the admin_peerEvents API, [^5] https://geth.ethereum.org/docs/interacting-with-geth/rpc/ns-admin which provides details such as node names, software, network IDs, compatible protocols, and IP addresses. This API alerts peer events to the subscriber in real-time, capturing both the addition and removal of peers, including those that connect/disconnect briefly. This information helps us calculate the number of active peer connections at each time, and better infer peer stability. For IP addresses, we call the ipinfo.io API [^6] https://ipinfo.io/developers/responses to obtain additional information such as Internet Service Provider (ISP) or Autonomous System (AS), hostname, timezone, region, city/country, and registered location. Besides regular full nodes, we also use our archive node to access the full state of the blockchain. In particular, we 1) retrieve the account balance/nonce at specific times (§5.3) and 2) fork the archive node to simulate the transaction validation process (§8.1). To accurately assess the attack cost amid price fluctuations, we obtain historical Ethereum/USD exchange rates through the Coingecko API. [^7] https://www.coingecko.com/api/documentation
@@
We modify Geth to store all messages received on the P2P network layer between September 1st, 2023 and January 25th, 2024. We deploy two nodes in the network to 1) avoid a single point of failure and 2) increase the coverage and robustness of our estimates. Those monitoring nodes enable us to observe transactions before they enter the txpool as described in Figure 2; modified components are highlighted in yellow, whereas the standard Geth components are in grey. We have integrated a custom message dispatcher that filters and timestamps specific transaction propagation messages (e.g., 0x02 and 0x08 messages). These messages are subsequently forwarded to a message queue, then batch-processed by a goroutine worker for database entry. This configuration ensures there is no delay between the capture of a message and its logging in the database. For each transaction hash, we only record the first message to reduce the size of the dataset. We store the type of messages, the content of the transactions, the origin node IDs, and the timestamp. This is critical for 1) estimating the active peer connections (§6.1), and 2) identifying spamming behavior and modified nodes within the P2P network (§6.2). Our two Geth nodes operate on Ubuntu 20.04.2 LTS systems, one with an AMD Ryzen Threadripper 3990X (64-core, 2.9 GHz) and the other with a 13th Gen Intel(R) Core(TM) i9-13900KF. Both machines have 256 GB and 64 GB of RAM, respectively, and are supported by NVMe SSDs. We raise the limit of peer connections in both Geth clients to 1,000. Both nodes are located in Europe and are operated independently. We keep the peering default configuration (no change in whitelist/blacklist). In total, we capture about 2.5 billion (2,493,695,017) unique 0x02 and 0x08 messages from 36,815 peers in the aforementioned timeframe.
3.3 Ethics
Following Tang et al. [48] on ethics of measuring blockchain P2P networks, our work consists of passive measurements (i.e., not submitting transactions). We do collect the Ethereum node IP addresses, but this is already publicly available information. As a part of responsible disclosure, we contacted the services that we discovered are running modified nodes and could be victims of this attack.
4 Modeling EDoS attack
We turn to formally modeling the Economic Denial of Sustainability (EDoS) attack in detail.
@@ 4.1 Threat model
We first formalize an amplified EDoS attack on the P2P blockchain network (Figure 1). The attack leads modified nodes (i.e., victims) to forward invalid transactions, and increases their outgoing traffic usage, thereby causing financial loss. The attacker attempts 1) to cause economic damage to modified nodes, 2) to degrade the quality of the service provided by modified nodes, and 3) to disrupt high-frequency MEV bots or block builders that listen to the services provided by modified nodes. Beyond blockchain ecosystem participants, cloud service providers also potentially have an incentive to support attacks on modified nodes to increase traffic usage and, consequently, revenue.
We next delineate the attack process. Let $\mathcal{A}$ denote the attacker (node), $\mathcal{M}$ be a modified node, and $\mathcal{R}$ be a regular (unmodified) node. $\mathcal{A}$ follows the procedure in Figure 3 to produce an invalid transaction, $tx_{invalid}$, of size $a$ bytes.
@@
    If the target $\mathcal{M}$ does not check the account balance, $\mathcal{A}$ can generate different addresses and send an infinite number of invalid transactions. The transaction cost is zero since there is no chance of inclusion.
@@
    If $\mathcal{M}$ checks the balance but not the nonce, $\mathcal{A}$ can send transactions with a previously used nonce, which also costs nothing. One necessary condition is for $\mathcal{A}$ to rely on an account with a long transaction history (i.e., having a high nonce allows $\mathcal{A}$ to reuse many past nonces).
@@
    If $\mathcal{M}$ checks balance/nonce but does not use the gas bump rule (§2), $\mathcal{A}$ can replicate transactions by altering a parameter (e.g., a data field), while maintaining the same sender, nonce, and gas price. $\mathcal{A}$ could incur transaction costs if the original (or one of the duplicated ones) ends up on the chain.
In the first two cases, $\mathcal{A}$ can specify an unreasonably high gas price to prioritize their transactions, without additional costs. $\mathcal{A}$ can also combine multiple strategies (e.g., duplicating many invalid transactions using a past nonce).
Next, $\mathcal{A}$ selects its target $\mathcal{M}$. If $\mathcal{A}$ receives any invalid transactions from a given peer, that peer is most likely $\mathcal{M}$. Alternatively, $\mathcal{A}$ can monitor various centralized services’ txpools, identify the service(s) that accept(s) invalid transactions, and send $tx_{invalid}$ through RPC endpoints. $\mathcal{A}$ finally sends $tx_{invalid}$ to $\mathcal{M}$, which accepts/inserts $tx_{invalid}$ to its txpool and forwards it with the rest of the network, including other $\mathcal{M}$ nodes and $\mathcal{R}$ nodes.
When all $\mathcal{M}$ nodes are densely connected and run by the same entity (as empirically observed in §6.2), $tx_{invalid}$ reaches all $\mathcal{M}$ nodes. Even if some $\mathcal{M}$ nodes belong to multiple entities, $tx_{invalid}$ would reach out to the rest as long as one of $\mathcal{M}$ from each entity connects.
Regular $\mathcal{R}$ nodes keep receiving $tx_{invalid}$ as a new transaction because $tx_{invalid}$ never gets into their txpool.

### v2

# Blockchain Amplification Attack

@@ Abstract
Strategies related to the blockchain concept of Extractable Value (MEV/BEV), such as arbitrage, front- or back-running create an economic incentive for network nodes to reduce latency. A modified node, that minimizes transaction validation time and neglects to filter invalid transactions in the Ethereum P2P network, introduces a novel attack vector—Blockchain Amplification Attack. An attacker exploits those modified nodes to amplify an invalid transaction thousands of times, posing a threat to the entire network. To illustrate attack feasibility and practicality in the current mainnet, we 1) identify thousands of similar attacks in the wild, 2) mathematically model propagation mechanism, 3) empirically measure model parameters from our two monitoring nodes, and 4) compare performance with existing Denial-of-Service attacks through local simulation. We show that an attacker can amplify network traffic at modified nodes by a factor of 3,600, and cause economic damages 13,800 times greater than the amount needed to carry out the attack. Despite these risks, aggressive latency reduction may still be profitable enough to justify the existence of modified nodes. To assess this trade-off, we 1) simulate the transaction validation process in the local network and 2) empirically measure the latency reduction by deploying our modified node in the Ethereum testnet. We conclude with a cost-benefit analysis of skipping validation and provide mitigation strategies against this attack.
@@ 1 Introduction
The journey of transactions in blockchain P2P networks en route to their ultimate validation remains understudied. The specifics of handling pending transactions vary according to the client’s software version and its unique implementation. Under the concept of Extractable Value (MEV/BEV) [9, 45]—obtaining extra profits by reordering transactions, centralized business entities have the incentive to reduce latency to deliver pending transactions. Being the fastest in this space (as little as 1 millisecond) attracts users, bots, and validators who profit from arbitrage or front/back running. This scenario bears a resemblance to the traditional stock market where trading firms physically shortened the network paths to stock exchange matching engines and optimized networking configuration for quicker trade execution [34]. Recently, we notice certain modified nodes tailor their configurations, sidestep transaction validation processes to shorten latency, and introduce invalid transactions that are not supposed to persist in the network.
@@
Figure 1: Overview of the lockchain Amplification Attack, illustrating how one invalid transaction amplifies in the network.
@@
We formalize the Blockchain Amplification Attack. In Figure 1, the adversary sends an invalid transaction (a red arrow) to the modified node(s) that collectively propagate an invalid transaction (blue arrows) to the rest of the network. In this setting, the attacker not only disables the modified nodes and causes economic damages through excessive traffic, but also spreads invalid transactions to slow down the entire network. This paper fully describes such potential attack opportunities, documents the prevalence of similar attacks in the wild, empirically/simulatively models the network damage, and quantifies the trade-offs associated with skipping validations. Due to the absence of penalty mechanism in the current P2P network, we demonstrate that the attack is practical, cost-effective, and a threat to the whole network.
@@
We first mathematically model the Blockchain Amplification Attack and pinpoint the parameters that affect how an attacker’s invalid transactions can get amplified in the whole network (i.e., amplification factor). While considering the specifics of each software/version on transaction forwarding, our model establishes a framework to quantify the increase in both the amount and the economic cost of egress traffic at the modified nodes.
@@
We next demonstrate the attack feasibility in practice. In particular, we analyze the set of pending transactions (called “txpool”) for centralized services. Some of them propagate transactions much faster than the others but deliver invalid transactions, suggesting the lack of validation checks. We identify 2,591 similar attack instances across 345 Ethereum blockchain addresses that target these services. We classify those instances based on the attacker’s observed strategies and estimate the attack cost, size, and intensity.
@@
To correctly estimate amplification factors (i.e., the attack impact) in the current mainnet, we design two customized monitoring Geth nodes (the most popular Ethereum node client) to infer the network topology. Our custom nodes scan the activities on the P2P layer and record specific types of messages before they undergo processing, resulting 2.5 billion observations over 5 months. This infrastructure allows us to propose a new method for inferring the number of active peer connections, which serves as a cost-effective/ethical alternative to previously proposed methods [13, 10, 4, 35]. We further show that 1.5% of nodes in our dataset exhibit lenient transaction validation (“modified nodes”). Many of those nodes appear to come from the same entity given the identical git commit in a node name, which we do not find in any public software repository. We show that the attacker can amplify outgoing network traffic at modified nodes by a factor of at least 3,600, and cause economic damages 13,800 times greater than the amount needed to carry out the attack.
@@
In addition, we conduct attack simulations on the local P2P network and confirm that our proposed attack can evict as many honest transactions (from both the txpool/the block) as the existing DoS attacks but at significantly lower costs.
@@
Furthermore, we experiment to quantify the benefits of skipping transaction validation, specifically the amount of time saved for each validation process. We first fork our archive node (the node stores all the historical states of the blockchain) to simulate/dissect the process of Geth’s transaction validation in the locally controlled environment. We find that validating one transaction takes roughly 1 millisecond and checking account status (nonce/balance) is the most time-consuming (86%). Moreover, we empirically measure the latency reduction by deploying the modified/regular nodes in the Ethereum testnet for 14 weeks.
@@
  • We mathematically define the Blockchain Amplification Attack and show the attack feasibility based on similar attacks in the wild.
  • By designing and deploying our custom monitoring nodes, we infer the Ethereum P2P network topology and illustrate the attack efficiency.
@@ 2.1 Blockchain and P2P network
Beyond protocol-level verifications, each software incorporates supplementary checks to mitigate DoS (Denial-of-Service) attacks. For example, Geth employs the “gas bump rule,” requiring the gas fee for a transaction with the same sender and nonce to be higher than the original transaction with a specified increase (e.g., 10% in Geth). The node should not accept replicated transactions with the same address, nonce, and gas price. Li et al. [35, §5.1] or Yaish et al.  [55, §4.1] provide detailed explanations.
@@
When users fetch the blockchain data or initiate a transaction, they typically interact with a P2P node through Remote Procedure Call (RPC) requests, and nodes execute these requests. Various publicly accessible RPC endpoints (e.g., Infura) allow users (rate limited) free access (including pending transactions in the txpool). Txpool view providers (e.g., bloXroute, Eden, Chainbound), in particular, offer a “better” view of txpool in terms of latency. For example, bloXroute employs a Blockchain Distribution Network (BDN) infrastructure to improve blockchain scalability [32, 3]. The globally distributed nodes index transactions to reduce the size, internally propagate without validations (“cut-through routing”), and employ dynamic routing to overcome network overhead.
@@ 2.2 Security in P2P network
Our proposed attack stems from not only DoS attack [33] but also Economic Denial of Sustainability (EDoS) attack. Hoff [29] first presented the EDoS concept in a 2008 blog post. EDoS, akin to a traditional Denial of Service (DoS) attack, focuses on inflicting financial losses by targeting the victim’s traffic usage [53]. Attackers exploit the knowledge of a victim server hosted on a cloud service by manipulating traffic usage to inflate bills, making the victim’s operation economically unsustainable. The same issue applies to blockchain P2P networks where many nodes have public IP addresses, and use cloud platforms such as AWS.
@@ 3.2 Private monitoring node data
To infer the network topology and estimate the attack impact, we modify Geth to store all messages received on the P2P network layer between September 1st, 2023 and January 25th, 2024. We deploy two nodes in the network to 1) avoid a single point of failure and 2) increase the coverage and robustness of our estimates. Those monitoring nodes enable us to observe transactions before they enter the txpool as described in Figure 2; modified components are highlighted in yellow, whereas the standard Geth components are in grey. We have integrated a custom message dispatcher that filters and timestamps specific transaction propagation messages (e.g., 0x02 and 0x08 messages). These messages are subsequently forwarded to a message queue, then batch-processed by a goroutine worker for database entry. This configuration ensures there is no delay between the capture of a message and its logging in the database. For each transaction hash, we only record the first message to reduce the size of the dataset. We store the type of messages, the content of the transactions, the origin node IDs, and the timestamp. This is critical for 1) estimating the active peer connections (§6.1), and 2) identifying spamming behavior and modified nodes within the P2P network (§6.2). Our two Geth nodes operate on Ubuntu 20.04.2 LTS systems, one with an AMD Ryzen Threadripper 3990X (64-core, 2.9 GHz) and the other with a 13th Gen Intel(R) Core(TM) i9-13900KF. Both machines have 256 GB and 64 GB of RAM, respectively, and are supported by NVMe SSDs. We raise the limit of peer connections in both Geth clients to 1,000. Both nodes are located in Europe and are operated independently. We keep the peering default configuration (no change in whitelist/blacklist). In total, we capture about 2.5 billion (2,493,695,017) unique 0x02 and 0x08 messages from 36,815 peers in the aforementioned timeframe.
To further profile each node within a network, we gather peer information from our nodes. This is done by subscribing to the admin_peerEvents API, [^5] https://geth.ethereum.org/docs/interacting-with-geth/rpc/ns-admin which provides details such as node names, software, network IDs, compatible protocols, and IP addresses. This API alerts peer events to the subscriber in real-time, capturing both the addition and removal of peers, including those that connect/disconnect briefly. This information helps us calculate the number of active peer connections at each time, and better infer peer stability. For IP addresses, we call the ipinfo.io API [^6] https://ipinfo.io/developers/responses to obtain additional information such as Internet Service Provider (ISP) or Autonomous System (AS), hostname, timezone, region, city/country, and registered location. Besides regular full nodes, we also use our archive node to access the full state of the blockchain. In particular, we 1) retrieve the account balance/nonce at specific times (§5.3) and 2) fork the archive node to simulate the transaction validation process (§8.1). To accurately assess the attack cost amid price fluctuations, we obtain historical Ethereum/USD exchange rates through the Coingecko API. [^7] https://www.coingecko.com/api/documentation
4 Modeling amplification
We turn to formally modeling the Blockchain Amplification Attack in detail.
@@ 4.1 Threat model
We first formalize an amplification attack on the P2P blockchain network (Figure 1). The attack leads modified nodes (i.e., victims) to forward invalid transactions, and increases their outgoing traffic usage and cost. The attacker attempts 1) to degrade the quality of the service provided by modified nodes, 2) to cause economic damage to modified nodes, and 3) to disrupt all nodes and users (especially high-frequency MEV bots or block builders that listen to the services provided by modified nodes). Beyond blockchain ecosystem participants, cloud service providers also potentially have an incentive to support attacks on modified nodes to increase traffic usage and, consequently, revenue.
We next delineate the attack process. The attacker follows the procedure in Figure 3 to produce an invalid transaction, $tx_{invalid}$, of size $a$ bytes.
@@
    If the target modified does not check the account balance, the attacker can generate different addresses and send an infinite number of invalid transactions. The transaction cost is zero since there is no chance of inclusion.
@@
    If the modified node checks the balance but not the nonce, the attacker can send transactions with a previously used nonce, which also costs nothing. One necessary condition is for the attacker to rely on an account with a long transaction history (i.e., having a high nonce allows the attacker to reuse many past nonces).
@@
    If the modified node checks balance/nonce but does not use the gas bump rule (§2), the attacker can replicate transactions while maintaining the same sender, nonce, and gas price. The attacker slightly alters one parameter (e.g., a data field) and re-signs the transaction to yield a different hash. The attacker could incur transaction costs if the original (or one of the duplicated ones) ends up on the chain.
In the first two cases, the attacker can specify an unreasonably high gas price to prioritize their transactions, without additional costs. The attacker can also combine multiple strategies (e.g., duplicating many invalid transactions using a past nonce).
Next, the attacker selects its target modified node. If the attacker receives any invalid transactions from a given peer, that peer is most likely the modified node. Alternatively, the attacker can monitor various centralized services’ txpools, identify the service(s) that accept(s) invalid transactions, and send $tx_{invalid}$ through RPC endpoints. The attacker finally sends $tx_{invalid}$ to one modified node, which accepts/inserts $tx_{invalid}$ to its txpool and forwards it with the rest of the network, including other modified nodes and regular nodes.
The modified nodes are tyipcally run by the same entity (as empirically observed in §6.2), they are most likely connected to minimize latencies. $tx_{invalid}$ would reach all the modified nodes. Even if some modified nodes belong to multiple entities, $tx_{invalid}$ would reach out to the rest as long as one of the nodes from each entity connects.
Regular nodes keep receiving $tx_{invalid}$ as a new transaction because $tx_{invalid}$ never gets into their txpool. All nodes consume an incoming traffic and CPU resources for transaction verification, illustrating the attack severity.